xx ReflectionSymmetry

ReflectionSymmetry

Erik Schnetter <schnetter@aei.mpg.de>

June 01 2004

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11avoidxoriginxupperxx 10Scope: privatex01  BOOLEAN

11Description: Stagger about the origin on the upper x boundary?

11                10  x01  Default: yes

11

11avoidxoriginxupperxy 10Scope: privatex01  BOOLEAN

11Description: Stagger about the origin on the upper y boundary?

11                10  x01  Default: yes

11

11avoidxoriginxupperxz 10Scope: privatex01  BOOLEAN

11Description: Stagger about the origin on the upper z boundary?

11                10  x01  Default: yes

11

11avoidxoriginxx 10Scope: privatex01  BOOLEAN

11Description: Stagger about the origin on the lower x boundary?

11                10  x01  Default: yes

11

11avoidxoriginxy 10Scope: privatex01  BOOLEAN

11Description: Stagger about the origin on the lower y boundary?

11                10  x01  Default: yes

11

11avoidxoriginxz 10Scope: privatex01  BOOLEAN

11Description: Stagger about the origin on the lower z boundary?

11                10  x01  Default: yes

11

11reflectionxupperxx10Scope: privatex01  BOOLEAN

11Description: Reflection symmetry at the upper x boundary

11                10  x01  Default: no

11

11reflectionxupperxy10Scope: privatex01  BOOLEAN

11Description: Reflection symmetry at the upper y boundary

11                10  x01  Default: no

11

11reflectionxupperxz10Scope: privatex01  BOOLEAN

11Description: Reflection symmetry at the upper z boundary

11                10  x01  Default: no

11

11reflectionxx  10Scope: privatex01  BOOLEAN

11Description: Reflection symmetry at the lower x boundary

11                10  x01  Default: no

11

11reflectionxy  10Scope: privatex01  BOOLEAN

11Description: Reflection symmetry at the lower y boundary

11                10  x01  Default: no

11

11reflectionxz  10Scope: privatex01  BOOLEAN

11Description: Reflection symmetry at the lower z boundary

11                10  x01  Default: no

11verbose  10Scope: privatex01  BOOLEAN

11Description: Produce screen output while applying boundary conditions

11                10  x01  Default: no

11

7 Interfaces

General Implements: reflectionsymmetry

8 Schedule This section lists all the variables which are assigned storage by thorn CactusNumerical/ReflectionSymmetry. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

SymmetryRegister x reflectionsymmetryxregister x register reflection symmetry boundaries

11x11Language:11c
11 x 11Type: 11function

11

BoundaryConditions x reflectionsymmetryxapply x apply reflection symmetries x

11x11Language:11c
11 x 11Type: 11function